WVU Corporation has a sustainable growth rate of 12.5%, total assets to sales ratio of 1.5, profit margin of 10%, and dividend payout of 50%. What is the firm’s target D/E ratio?
Answer #1

Sustainable growth Rate = ROE(1 - Dividend Payout)

=> 0.125 = ROE(1 - 0.50)

=> ROE = 0.25

=> Net Income / Total Equity = 0.25

Profit Margin = Net Income / Sales = 0.10 ... (I)

Total Assets / Sales = 1.5 ... (II)

(II)/ (I) => (Total Assets / Sales) / (Net Income / Sales) = 1.5/ 0.10

=> Total Assets / Net Income = 15

=> Net Income = Total Assets / 15 ... (III)

Net Income / Total Equity = 0.25

Substituting (III)

=> (Total Assets / 15) / Total Equity = 0.25

=> Total Assets / Total Equity = 0.25*15 = 3.75

Assets = Debt + Equity , Substituting in above equation,

(Total Debt + Total Equity) / Total Equity = 3.75

=> 1 + Total Debt/Total Equity = 3.25

=> Total Debt/Total Equity = 2.75